Evaluate the triple integral of f(x,y,z)=z(x^2+y^2+z^2)^(−3/2) over the part of the ball x^2+y^2+z^2≤16 defined by z≥2.
1
Expert's answer
2020-11-18T17:26:47-0500

Converting the triple integrals to spherical coordinates, we have;

"f(ρ,ϕ,θ)=ρcos(ϕ)ρ^{−3/2}\\

f(ρ,ϕ,θ)=cos(ϕ)ρ^{−1/2}"


Then for the bounds:

"ρ^2≤16\\

ρ≤4"


and;

"ρcos(ϕ)≥2\\

\dfrac{2}{cos(ϕ)}≤ρ"



So:

"\dfrac{2}{cos(ϕ)}≤ρ≤4"


Then since the plane crosses the sphere at z=2 where ρ=4, and z = ρcos(ϕ):


"4cos(ϕ)=2\\

ϕ=π/3"


So:

"0\u2264\u03d5\u2264\u03c0\/3"


Since it is a sphere I have: "0\u2264\u03b8\u22642\u03c0"


So my bounds are:

"\frac{2}{cos(ϕ)}≤ρ≤4\\
0≤θ≤2π"

Over:

"f(\u03c1,\u03d5,\u03b8)=cos(\u03d5)\u03c1^{\u22121\/2}"


With "dV=\u03c1^2sin(\u03d5)d\u03c1d\u03d5d\u03b8",


The integral can be written as:


"∫_0^{2π}∫_0^{π/3}∫_{\frac{2}{cos(ϕ)}}^4ρ^{−1/2}cos(ϕ)ρ^2sin(ϕ)dρdϕdθ=\\"

"∫_0^{2π}∫_0^{π/3}∫_{\frac{2}{cos(ϕ)}}^4ρ^{-3/2}cos(ϕ)sin(ϕ)dρdϕdθ="


"∫_0^{2π}dθ∫_0^{π/3}cos(ϕ)sin(ϕ)dϕ∫_{\frac{2}{cos(ϕ)}}^4ρ^{−3/2}dρ="

"2π× \frac{√3}{4}×∫_{\frac{2}{cos(ϕ)}}^4ρ^{−3/2}dρ="


"\frac{π√3}{2} × \frac{2}{√3} = π"